The sum of 1111 1 1111 in binary equals

1) 11111

2) 0000

3) 2222

4) 11110
Answer #1

  • (1111)2 = 1510 (1)2 =110
  • (1111)2 + 12 + (1111)2 = (15+1+15)10 = (31)10 =(11111)2